Kim and Kanye have been dating for years and are now thinking about getting married. As a financially sophisticated couple, they want to think through the tax implications of their potential union.

Suppose Kim and Kanye both earn $70,000 (so their combined income is $140,000). Using the tax bracket information in Table 1.2 (or the Excel file available on MyFinanceLab with the same information), calculate the combined tax bill that they would pay if they remain single, and compare that to taxes they would pay if they were married and filed a joint return.

Now suppose that Kim and Kanye both earn $100,000 (so their combined income is $200,000). Calculate the combined tax bill that they would pay if they remain single, and compare that to the taxes that they would pay if they were married and filed a joint return.

What differences do you find in parts (a) and (b)? What is the cause of these differences?

Table 1.2

                          Taxable Income

Tax Rates

Individual Return

Joint Returns

10.0%

$0 to $9,075

$0 to $18,150

15.0%

$9,076 to $36,900

$18,151 to $73,800

25.0%

$36,901 to $89,350

$73,801 to $148,850

28.0%

$89,351 to $186,350

$148,851 to $226,850

33.0%

$186,351 to $405,100

$226,851 to $405,100

35.0%

$405,101 to $406,750

$405,101 to $457,600

39.6%

Over $406,750

Over $457,600
Answer #1

(a) Individual Incomes = $70000, Total Income = $140000

If they pay taxes individually,
Tax liability for each = 9075 * 10% + (36900 - 9075) * 15% + (70000 - 36900) * 25% = $13356.25
Total Paid by them = 13356.25 * 2 = $26712.5

If they pay joint taxes,
Total Tax liability = 18150 * 10% + (73800 - 18150) * 15% + (140000 - 73800) * 25% = $26712.5

(b)

Individual Incomes = $100000, Total Income = $200000

If they pay taxes individually,
Tax liability for each = 9075 * 10% + (36900 - 9075) * 15% + (89350 - 36900) * 25% + (100000 - 89350) * 28% = $21175.75
Total Paid by them = 21175.75 * 2 = $42351.5

If they pay joint taxes,
Total Tax liability = 18150 * 10% + (73800 - 18150) * 15% + (148850 - 73800) * 25% + (200000 - 148850) * 28% = $43247

In part (a), the tax liability remains same whether they pay the taxes individually or together. However, in case (b) the tax liabilities increases if they pay taxes together.

This happens because the tax brackets do not double when the income increases. For any individual income under $74425 ($148,850/2), the tax liabilities would remain the same. However if the individual income increases above $74,425, the tax buckets change for individual returns and collective returns.

For example, if the individual income is $80,000, which is greater than $74,425 but less than $89,350. In this case if the taxes are filed individually, the highest tax bracket will be 25%. But if joint returns are filed the highest tax bracket will be 28% (80000*2 = 160000, which lies in 28% bucket).